# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 5.3 Randomized Lower-Bound for $\delta$: Distance between vertices and facets
as desired. ∎
5.4 Randomized Lower-Bound for $r$: Inner Radius of a Ridge Projected onto $(d-1)$-Dimensional Space
In this section, we demonstrate that for any ridge $R$ of the polytope $P$, wherein $R\cap W$ is a vertex of $P\cap W$, its inner radius—after projection onto the subspace orthogonal to the adjacent edge of $P\cap W$—is at least $\Omega(d^{-4}L^{-2})$ with high probability. Essentially, this establishes that the parameter $r$, as referred to in Lemma 36, is at least $\Omega(d^{-4}L^{-2})$ with good probability. We remark that Lemma 42 has no analogue when $d=2$ and will require more technical effort. Its proof is similar to Lemma 4.1.1 (Distance bound) in [ST04].
Lemma 42 (Randomized Lower-bound for $r$).
@@ 5.4 Randomized Lower-Bound for $r$: Inner Radius of a Ridge Projected onto $(d-1)$-Dimensional Space
Proof.
We may assume without loss of generality that $I=[d]$ and $J=[d-1]$. Apply the change of variables $\phi$ as in Definition 18 to $\{a_{i}:i\in[d]\}$ and obtain
@@
where $\operatorname{vol}_{d-1}(\cdot)$ is the volume function of $(d-1)$-dimensional simplex, $\bar{\mu}_{i}(t,\theta,b_{i})=\mu_{i}(t\theta+R_{\theta}(b_{i}))$ is the induced probability density of $b_{i}$, which is $L$-log-Lipschitz, and $1[\cdot]$ denotes the indicator function. Write $S$ for the event that
@@
With probability $1$ over the choice of $\theta,t,a_{d+1},\dots,a_{n}$, the inner term satisfies all the conditions of Lemma 45. Specifically, since the value of $1[E_{[d]}]$ is already fixed, the intersection $(t\theta+\theta^{\perp})\cap W$ is a line. Let $\ell\subseteq\mathbb{R}^{d-1}$ be the image of such line under the inverse change of variables $\phi^{-1}$, i.e. $(t\theta+\theta^{\perp})\cap W=t\theta+R_{\theta}(\ell)$. Then the event $A_{[d-1]}$ is equivalent to $\ell\cap\operatorname{conv}(b_{1},\dots,b_{d-1})\neq\emptyset$. From Lemma 19, the joint probability distribution of $b_{1},\dots,b_{d}$ is thus proportional to
@@
Applying Lemma 45 to the term $\Pr_{b_{1},\dots,b_{d}}[S]$ we find
@@
using Fubini’s theorem for the equality and the fact that $\neg D^{\prime}$ implies $\neg D$ for the final inequality. ∎
Lemma 43.
Let $\ell\subseteq\mathbb{R}^{k}$ be an affine line and $x\in\mathbb{R}^{k}$ be point. Suppose $w\in\mathbb{S}^{k-1}\setminus\{0\}$ points in the direction of $\ell$, i.e., that $\ell+w=\ell$. Then $\operatorname{dist}(x,\ell)=\operatorname{dist}(\pi_{w^{\perp}}(x),\pi_{w^{\perp}}(\ell))$.
Proof.
Let $\{z\}=\ell\cap(x+w^{\perp})$. This intersection is non-empty because $\dim(\ell)+\dim(x+w^{\perp})=k$, and is a singleton because if $z,z^{\prime}$ were two distinct points in this set then $z-z^{\prime}\in w\mathbb{R}\cap w^{\perp}=\{0\}$.
@@
Lemma 44.
Let $H\subseteq\mathbb{R}^{k}$ be a co-dimension $1$ hyperplane. Let $p_{1},\dots,p_{k}\in H$ and $p_{k+1}\in\mathbb{R}^{k}$, and assume $\lambda_{1},\dots,\lambda_{k+1}\geq 0$ are such that $\sum_{i=1}^{k+1}\lambda_{i}=1$.
@@
using that $y^{\top}p_{i}=t$ for all $i=1,\dots,k$. ∎
Lemma 45 (Randomized lower bound for $r$ after change of variables).
where $\bar{\mu}_{i}$ is $L$-log-Lipschitz for each $i\in[d]$. Let $D^{\prime}$ denote the event that the set $\{b_{1},\dots,b_{d}\}$ has Euclidean diameter of at most $3$. Given any fixed one-dimensional line $\ell\subseteq\mathbb{R}^{d-1}$, we have that
Proof.
We can bound the distance from $\ell$ to $\partial\operatorname{conv}(b_{1},\dots,b_{d-1})$.
@@
$=$ $\operatorname{dist}\big{(}\pi_{w^{\perp}}(\ell),\pi_{w^{\perp}}(\partial\operatorname{conv}(b_{1},\dots,b_{d-1}))\big{)}$ (By Lemma 43)
@@
$=$ $\min_{i\in[d-1]}\lambda_{i}\cdot\operatorname{dist}(\pi_{w^{\perp}}(b_{i}),\operatorname{affhull}(\pi_{w^{\perp}}(b_{j}):j\in[d-1],j\neq i))$ (By Lemma 44)
@@
where in the fifth step, $\lambda\in\mathbb{R}^{d-1}_{\geq 0}$ is the unique solution to $\sum_{i=1}^{d-1}\lambda_{i}b_{i}=\ell\cap\operatorname{conv}(b_{1},\ldots,b_{d-1})$ and $\sum_{i=1}^{d-1}\lambda_{i}=1$. Additionally, assume $\ell=w\mathbb{R}$ for a non-zero $w\in\mathbb{S}^{d-2}$. Abbreviate, for $k\in[d-1]$,
@@
Let $T$ denote the event that $\ell\cap\operatorname{conv}(b_{1},\dots,b_{d-1})\neq\emptyset$. We now find using the union bound, for any $\alpha,\beta>0$,
@@
By Lemma 46 we know that $\Pr[\min_{i\in[d-1]}\lambda_{i}<\alpha\mid D^{\prime}\wedge T]\leq 0.05$ for $\alpha=\frac{1}{120d^{2}L}$. By Lemma 47, we know that $\Pr[\min_{k\in[d-1]}r_{k}<\beta\mid T]\leq\sum_{k\in[d-1]}\Pr[r_{k}<\beta\mid T]\leq 0.05$ for $\beta=\frac{1}{160d^{2}L}$. This proves the lemma. ∎
The following two Lemmas are the key to prove Lemma 45. Recall from the proof of Lemma 45 that $\lambda\in\mathbb{R}^{d-1}_{\geq 0}$ is the unique solution to $\sum_{i=1}^{d-1}\lambda_{i}b_{i}=\ell\cap\operatorname{conv}(b_{1},\ldots,b_{d-1})$ and $\sum_{i=1}^{d-1}\lambda_{i}=1$. First, we use Lemma 46 to show that the every convex parameter $\lambda_{i}$ is at least $\Omega(1/d^{2}L)$ with constant probability.
Lemma 46 (Lower-bound for Convex Parameters of Vertices on the Ridge).
where each $\bar{\mu}_{i}:\mathbb{R}^{d-1}\to\mathbb{R}_{+}$ is $L$-log-Lipschitz. Given any one-dimensional line $\ell\subseteq\mathbb{R}^{d-1}$ and conditional on $\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])\neq\emptyset$. Let $\lambda\in\mathbb{R}^{d-1}_{+}$ be the unique solution to $\sum_{i=1}^{d-1}\lambda_{i}b_{i}\in\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])$. Let $D^{\prime}$ denote the event that $\forall i,j\in[d]$, $\|b_{i}-b_{j}\|\leq 3$. Then we have
Proof.
By using the union bound, it suffices to prove for each $i\in[d-1]$ that
@@
Fix any $i\in[d-1]$, without loss of generality $i=1$. We can assume $\ell=w\mathbb{R}$ for a non-zero $w\in\mathbb{S}^{d-2}$. Thus $\lambda$ is defined to satisfy $\sum_{j=1}^{d-1}\lambda_{j}\pi_{w^{\perp}}(b_{j})=0$.
For any given values of $b_{1}-b_{j}$, $j\in[d]$, which determine the shape of the simplex $\operatorname{conv}(b_{j}:j\in[d])$, we prove the result using the randomness in $\pi_{w^{\perp}}(b_{1})$, the position of the simplex in the space $w^{\perp}$. For the remainder of this proof, we can consider all $b_{j},j\in[d]$ to be functions of $b_{1}$. If we furthermore fix any value for $w^{\top}b_{1}$ then $\operatorname{vol}(\operatorname{conv}(b_{j}:j\in[d])$ is fixed, hence $\pi_{w^{\perp}}(b_{1})$ has probability density $\mu^{\prime}(\pi_{w^{\perp}}(b_{1}))\propto\prod_{j=1}^{d}\mu_{j}(b_{j})$, which is $dL$-log-Lipschitz in $\pi_{w^{\perp}}(b_{1})$ with respect to the $d-2$-dimensional Lebesgue measure on $w^{\perp}$.
Write $M=\operatorname{conv}(\pi_{w^{\perp}}(b_{1}-b_{j}):j\in[d-1])\subseteq w^{\perp}$, for which we can see that $\pi_{w^{\perp}}(b_{1})\in M$ if and only if $\lambda\geq 0$. It remains to show that
@@
For any $j\in[d-1]$, let $l_{j}:M\to[0,1]$ be the function sending any point to its $j$’th convex coefficient, i.e., the functions satisfy $\sum_{j=1}^{d-1}l_{j}(x)=1$ and $\sum_{j=1}^{d-1}l_{j}(x)\cdot\pi_{w^{\perp}}(b_{1}-b_{j})=x$ for every $x\in M$. For any $1\geq\alpha\geq 0$, observe that $l_{1}$ takes values in the interval $[\alpha,1]$ on the set $(1-\alpha)M$. Hence we get
@@
$\geq(1-\alpha)^{d-2}\max_{x\in M}e^{-dL\|\alpha x\|}.$ (By $d$-log-Lipschitzness of $\mu^{\prime}$)
@@
where the last line comes from $L\geq 1$. Thus (20) holds as desired. ∎
In the second part, we lower bound the distance between each vertex $b_{j}$ (where $j\in[d-1]$) and the $(d-3)$-dimensional hyperplane spanned by the other vertices $\operatorname{affhull}(b_{j}:j\in[d-1],j\neq i)$. Specifically, we show the following lemma:
Lemma 47.
where each $\bar{\mu}_{i}:\mathbb{R}^{d-1}\to[0,1]$ is $L$-log-Lipschitz. Given any one-dimensional line $\ell\subseteq\mathbb{R}^{d-1}$, and let $w\in\mathbb{S}^{d-2}$ be any unit direction of $\ell$. For any $i\in[d-1]$ we have
$\Pr\Big{[}\operatorname{dist}\left(\pi_{w^{\perp}}(b_{i}),\operatorname{affhull}(\pi_{w^{\perp}}(b_{j}):j\in[d-1],j\neq i)\right)\geq\frac{1}{160d^{2}L}~{}\Big{|}~{}$ $\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])\neq\emptyset\Big{]}\geq 1-\frac{1}{20d}.$
Proof.
In the following arguments, we condition on $\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])\neq\emptyset$. Write $\pi=\pi_{\ell^{\perp}}$ for the orthogonal projection onto $\ell^{\perp}$. Without loss of generality, set $i=d-1$ and assume that $\ell$ is a linear subspace, i.e., $\pi(\ell)=0$.
We start with a coordinate transformation. Let $\phi\in w^{\perp}\cap\mathbb{S}^{d-2}$ denote the unit vector satisfying $\phi^{\top}b_{1}=\phi^{\top}b_{j}>0$ for all $j=1,\dots,d-1$. Note that $\phi$ is uniquely defined almost surely: $w^{\perp}$ is a $(d-2)$-dimensional linear space and we impose $(d-3)$ linear constraints $\{\phi^{\top}a_{1}=\phi^{\top}a_{j},\forall j\in[d-2]\}$. Almost surely, these give an one-dimensional linear subspace which, after adding the unit norm and $b_{1}^{\top}\phi>0$ constraint, leaves a unique choice of $\phi$.
Now define $h\in\mathbb{R}$ by $h=\phi^{\top}b_{1}$ and define $\alpha\in\mathbb{R}$ by $\alpha h=-\phi^{\top}b_{d-1}$. Since $0\in\operatorname{conv}(\pi(b_{i}):i\in[d-1])$ but $\phi^{\top}b_{i}>0$ for all $i\in[d-2]$, we must have $\alpha\geq 0$ for otherwise $\phi$ would separate $\operatorname{conv}(\pi(b_{i}):i\in[d-1])$ from 00. Again from almost-sure non-degeneracy we get $\alpha>0$ and $h\neq 0$. We define the following coordinate transformation:
$b_{j}$ $=h\phi+c_{j},\quad\forall j\in[d-2]$
$b_{d-1}$ $=-\alpha h\phi+c_{d-1}$
where for each $j\in[d-1]$, $c_{j}\in\phi^{\perp}\cap\operatorname{affhull}(b_{1},\ldots,b_{d-1})$ has $(d-3)$ degrees of freedom. From here on out, we consider the vertices $(b_{1},\ldots,b_{d-1})$ to be a function of $(h,\alpha,\phi,c_{1},\dots,c_{d-1})$. Again by Lemma 19, the induced joint probability density on the random variables $(h,\alpha,\phi,c_{1},\dots,c_{d-1},b_{d})$, is proportional to
[…]
Condition on the exact values of $(\alpha,\phi,c_{1},\dots,c_{d-1},b_{d})$. Note that the event $0\in\operatorname{conv}(\pi(b_{i}):i\in[d-1])$ depends only on these variables and not on $h$, and the same is true for $\operatorname{vol}_{d-3}(\operatorname{conv}(\pi_{w^{\perp}}(c_{1}),\dots,\pi_{w^{\perp}}(c_{d-2})))$. By Lemma 19, the induced probability density on $h$ is now proportional to
Next, consider the volume term. We can write $\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\dots,b_{d}))$ as a constant depending on $d$ times the absolute value of the determinant of matrix
$\begin{bmatrix}(b_{1}-b_{d})^{\top}\\ \vdots\\ (b_{d-1}-b_{d})^{\top}\end{bmatrix}=\begin{bmatrix}(h\phi+c_{1}-b_{d})^{\top}\\ \vdots\\ (h\phi+c_{d-2}-b_{d})^{\top}\\ (-\alpha h\phi+c_{d-1}-b_{d})^{\top}\end{bmatrix}=\begin{bmatrix}(c_{1}-b_{d})^{\top}\\ \vdots\\ (c_{d-2}-b_{d})^{\top}\\ (c_{d-1}-b_{d})^{\top}\end{bmatrix}+h\cdot\begin{bmatrix}1\\ \vdots\\ 1\\ -\alpha\end{bmatrix}\phi^{\top}.$
where $\theta\in\mathbb{S}^{d-1}$ denotes a normal vector to $\operatorname{affhull}(b_{1},\ldots,b_{d-1})$. Define
$B:=\begin{bmatrix}(c_{1}-b_{d})^{\top}\\ \vdots\\ (c_{d-2}-b_{d})^{\top}\\ (c_{d-1}-b_{d})^{\top}\\ \theta^{\top}\end{bmatrix},v:=\begin{bmatrix}1\\ \vdots\\ 1\\ -\alpha\\ 0\end{bmatrix}.$
Then by the matrix determinant lemma, we can write the volume as the absolute value of an affine function of $h$ (which is a convex function):
$k(h):=\operatorname{vol}_{d-1}(\operatorname{conv}(b_{1},\ldots,b_{d}))$ $=\left|\det(B)+hv\phi^{\top}\right|$
$=\left|\det(B)(1+\phi^{\top}B^{-1}v\cdot h)\right|$
To finalize the argument, we write $\operatorname{dist}(\pi(b_{i}),\operatorname{affhull}(\pi(b_{j}):j\in[d-1],j\neq i))=|(1+\alpha)h|$. It follows that the signed distance $(1+\alpha)h$ has a probability density function proportional to the product of a $dL$-log-Lipschitz function and a convex function. The result follows from Lemma 48 by plugging in the signed distance $(1+\alpha)h$ and $K=dL,\varepsilon=\frac{1}{160d^{2}L}$. ∎
Lemma 48.

### v3

@@ 5.3 Randomized Lower-Bound for $\delta$: Distance between vertices and facets
as desired.
5.4 Randomized Lower-Bound for $r$: Inner Radius of a Ridge Projected onto (d1)(d-1)-Dimensional Subspace
In the next sections, we demonstrate that for any ridge $R$ of the polytope $P$, wherein $R\cap W$ is a vertex of $P\cap W$, its inner radius—after projection onto the subspace orthogonal to the adjacent edge of $P\cap W$—is at least $\Omega(d^{-4}L^{-2})$ with high probability. Essentially, this establishes that the parameter $r$, as referred to in Lemma 5.6, is at least $\Omega(d^{-4}L^{-2})$ with good probability. We remark that Lemma 5.16 does not have an analogue when $d=2$. Moreover, it will require substantially more technical effort. Its proof is similar to Lemma 4.1.1 (Distance bound) in [49], their main technical result. In an effort to help the ease of understanding the larger structure, some lemmas will be stated while their proofs will be given in later sections.
Lemma 5.16 (Randomized Lower-bound for $r$).
@@ 5.4 Randomized Lower-Bound for $r$: Inner Radius of a Ridge Projected onto (d1)(d-1)-Dimensional Subspace
Then Lemma 5.16 will quickly from its lower-dimensional equivalent:
Lemma 5.17 (Randomized lower bound for $r$ after change of variables).
where $\bar{\mu}_{i}$ is $L$-log-Lipschitz for each $i\in[d]$. Let $D^{\prime}$ denote the event that the set $\{b_{1},\dots,b_{d}\}$ has Euclidean diameter of at most $3$. Fox any fixed one-dimensional line $\ell\subseteq\mathbb{R}^{d-1}$, we have that
The proof of Lemma 5.17 will be presented in Section 5.5.
Proof 5.18 (Proof of Lemma 5.16).
We may assume without loss of generality that $I=[d]$ and J=[d1]J=[d-1]. Apply the change of variables $\phi$ as in Definition 2.12 to $\{a_{i}:i\in[d]\}$ and obtain
@@
where $\operatorname{vol}_{d-1}(\cdot)$ is the volume function of (d1)(d-1)-dimensional simplex in its spanning hyperplane, $\bar{\mu}_{i}(t,\theta,b_{i})=\mu_{i}(t\theta+R_{\theta}(b_{i}))$ is the induced probability density of $b_{i}$, which is $L$-log-Lipschitz, and $1[\cdot]$ denotes the indicator function. Write $S$ for the event that
@@
With probability $1$ over the choice of $\theta,t,a_{d+1},\dots,a_{n}$, the inner term satisfies all the conditions of Lemma 5.17. Specifically, since the value of $1[E_{[d]}]$ is already fixed, the intersection $(t\theta+\theta^{\perp})\cap W$ is a line. Let $\ell\subseteq\mathbb{R}^{d-1}$ be the image of such line under the inverse change of variables $\phi^{-1}$, i.e., $(t\theta+\theta^{\perp})\cap W=t\theta+R_{\theta}(\ell)$. Then the event $A_{[d-1]}$ is equivalent to $\ell\cap\operatorname{conv}(b_{1},\dots,b_{d-1})\neq\emptyset$. From Lemma 2.13, the joint probability distribution of $b_{1},\dots,b_{d}$ is thus proportional to
@@
Applying Lemma 5.17 to the term $\Pr_{b_{1},\dots,b_{d}}[S]$ we find
@@
using Fubini’s theorem for the equality and the fact that $\neg D^{\prime}$ implies $\neg D$ for the final inequality.
5.5 Proof of Lemma 5.17: Randomized lower bound for $r$ after change of variables
In this section, we deliver the proof of Lemma 5.17. We use the following two technical lemmas. As in from the proof of Lemma 5.17 we define $\lambda\in\mathbb{R}^{d-1}_{\geq 0}$ to be the unique solution to $\sum_{i=1}^{d-1}\lambda_{i}b_{i}=\ell\cap\operatorname{conv}(b_{1},\ldots,b_{d-1})$ and $\sum_{i=1}^{d-1}\lambda_{i}=1$. First, we describe Lemma 5.19 which we will use to show that every convex parameter $\lambda_{i}$ is at least $\Omega(1/d^{2}L)$ with constant probability.
Lemma 5.19 (Lower-bound for Convex Parameters of Vertices on the Ridge).
where each $\bar{\mu}_{i}:\mathbb{R}^{d-1}\to\mathbb{R}_{+}$ is $L$-log-Lipschitz. Let one-dimensional line $\ell\subseteq\mathbb{R}^{d-1}$ and conditional on $\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])\neq\emptyset$. Let $\lambda\in\mathbb{R}^{d-1}_{+}$ be the unique solution to $\sum_{i=1}^{d-1}\lambda_{i}b_{i}\in\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])$. Let $D^{\prime}$ denote the event that $\forall i,j\in[d]$, $\|b_{i}-b_{j}\|\leq 3$. Then we have
We will prove Lemma 5.19 in Section 5.6.
Secondly, we present a lemma to lower bound the distance between each vertex $b_{j}$ (where $j\in[d-1]$) and the (d3)(d-3)-dimensional hyperplane spanned by the other vertices $\operatorname{affhull}(b_{j}:j\in[d-1],j\neq i)$. Specifically, we show the following lemma:
Lemma 5.20.
where each $\bar{\mu}_{i}:\mathbb{R}^{d-1}\to[0,1]$ is $L$-log-Lipschitz. Given a one-dimensional line $\ell\subseteq\mathbb{R}^{d-1}$, let $w\in\mathbb{S}^{d-2}$ be any unit direction of $\ell$. For any $i\in[d-1]$ we have
$\Pr\Big[\operatorname{dist}\left(\pi_{w^{\perp}}(b_{i}),\operatorname{affhull}(\pi_{w^{\perp}}(b_{j}):j\in[d-1],j\neq i)\right)\geq\frac{1}{160d^{2}L}~\Big|$~ $\ell\cap\operatorname{conv}(b_{i}:i\in[d-1])\neq\emptyset\Big]$
$\geq 1-\frac{1}{20d}.$
We defer the proof of Lemma 5.20 to 5.27. In addition to the above two lemmas, we need the following two basic linear algebraic statements.
Fact 7.
Let $\ell\subseteq\mathbb{R}^{k}$ be an affine line and $x\in\mathbb{R}^{k}$ be a point. Suppose $w\in\mathbb{S}^{k-1}\setminus\{0\}$ points in the direction of $\ell$, i.e., that $\ell+w=\ell$. Then $\operatorname{dist}(x,\ell)=\operatorname{dist}(\pi_{w^{\perp}}(x),\pi_{w^{\perp}}(\ell))$.
Proof 5.21.
Let $\{z\}=\ell\cap(x+w^{\perp})$. This intersection is non-empty because $w$ points in the direction along $\ell$ and so must intersect any affine subspace orthogonal to $w$. The intersection is a singleton because if $z,z^{\prime}$ were two distinct points in this set then $z-z^{\prime}\in w\mathbb{R}\cap w^{\perp}=\{0\}$.
@@
Fact 8.
Let $H\subseteq\mathbb{R}^{k}$ be a hyperplane, and let $p_{1},\dots,p_{k}\in H$ and $p_{k+1}\in\mathbb{R}^{k}$, and assume $\lambda_{k+1}\geq 0$.
@@
using that $y^{\top}p_{i}=t$ for all $i=1,\dots,k$.
Assuming Lemma 5.19 and Lemma 5.20, we can prove Lemma 5.17.
Proof 5.23 (Proof of Lemma 5.17).
We can bound the distance from $\ell$ to $\partial\operatorname{conv}(b_{1},\dots,b_{d-1})$. Note that generically $w$ is not parallel to any direction in $\operatorname{affhull}(b_{1},\dots,b_{d-1})$ which makes it so that $\pi_{w^{\perp}}(\operatorname{conv}(b_{1},\dots,b_{d-1}))$ is a simplex of the same dimension as $\operatorname{conv}(b_{1},\dots,b_{d-1})$.
@@
$=$ $\operatorname{dist}\big(\pi_{w^{\perp}}(\ell),\pi_{w^{\perp}}(\partial\operatorname{conv}(b_{1},\dots,b_{d-1}))\big)$ (By Fact 7)
@@
$=$ $\min_{i\in[d-1]}\lambda_{i}\cdot\operatorname{dist}(\pi_{w^{\perp}}(b_{i}),\operatorname{affhull}(\pi_{w^{\perp}}(b_{j}):j\in[d-1],j\neq i))$ (By Fact 8)
@@
Where in the second step, we use the fact that $\pi_{w^{\perp}}$ is an affine isomorphism restricting to $b_{1},\dots,b_{d-1}$, thus taking the boundary of $\operatorname{conv}(\pi_{w^{\perp}}(b_{1}),\dots,\pi_{w^{\perp}}(b_{d-1}))$ commutes with the projection $\pi_{w}^{\perp}$. In the fifth step, $\lambda\in\mathbb{R}^{d-1}_{\geq 0}$ is the unique solution to $\sum_{i=1}^{d-1}\lambda_{i}b_{i}=\ell\cap\operatorname{conv}(b_{1},\ldots,b_{d-1})$ and $\sum_{i=1}^{d-1}\lambda_{i}=1$. Additionally, assume $\ell=w\mathbb{R}$ for a non-zero $w\in\mathbb{S}^{d-2}$. Abbreviate, for $k\in[d-1]$,
@@
Let $T$ denote the event that $\ell\cap\operatorname{conv}(b_{1},\dots,b_{d-1})\neq\emptyset$. We now find using a union bound, for any $\alpha,\beta>0$,
@@
By Lemma 5.19 we know that $\Pr[\min_{i\in[d-1]}\lambda_{i}<\alpha\mid D^{\prime}\wedge T]\leq 0.05$ for $\alpha=\frac{1}{120d^{2}L}$. By Lemma 5.20, we know that $\Pr[\min_{k\in[d-1]}r_{k}<\beta\mid T]\leq\sum_{k\in[d-1]}\Pr[r_{k}<\beta\mid T]\leq 0.05$ for $\beta=\frac{1}{160d^{2}L}$. This proves the lemma.
5.6 Proof of Lemma 5.19
Now we show that, with good probability, the convex multipliers $\lambda$ are not too small.
Proof 5.24 (Proof of Lemma 5.19).
By using a union bound, it suffices to prove for each $i\in[d-1]$ that
@@
Fix any $i\in[d-1]$, without loss of generality $i=1$. Recall that $w\in\mathbb{S}^{d-1}$ is such that $\ell=\ell+w$, and hence that $\pi_{w^{\perp}}(\ell)$ is a singleton point. For ease of exposition we assume that the plane $w^{\perp}$ is coordinatized such that $\pi_{w^{\perp}}(\ell)=0$ and hence $\ell=w\mathbb{R}$. Thus $\lambda$ is defined to satisfy $\sum_{j=1}^{d-1}\lambda_{j}\pi_{w^{\perp}}(b_{j})=0$.
Using the principle of deferred decision, we fix the values of $b_{1}-b_{j}$, $j\in[d]$. This determines the shape of the simplex $\operatorname{conv}(b_{j}:j\in[d])$, including its volume. The remainder of this proof will use the randomness in the position of the simplex in the subspace orthogonal to the line, which we represent using $\pi_{w^{\perp}}(b_{1})$. For the remainder of this proof, we can consider all $b_{j},j\in[d]$ to be functions of $b_{1}$. The position $\pi_{w^{\perp}}(b_{1})$ has probability density $\mu^{\prime}(\pi_{w^{\perp}}(b_{1}))\propto\prod_{j=1}^{d}\mu_{j}(b_{j})$, which is $dL$-log-Lipschitz in $\pi_{w^{\perp}}(b_{1})$ with respect to the (d2)(d-2)-dimensional Lebesgue measure on $w^{\perp}$.
Define $M=\operatorname{conv}(\pi_{w^{\perp}}(b_{1}-b_{j}):j\in[d-1])\subseteq w^{\perp}$ and note that, due to our fixing the values of $b_{1}-b_{j}$ in the previous paragraph, the shape $M$ is fixed and we can see that that $\pi_{w^{\perp}}(b_{1})\in M$ if and only if $\lambda\geq 0$. It remains to show that
@@
For any $j\in[d-1]$, let $l_{j}:M\to[0,1]$ be the function sending any point to its $j$’th convex coefficient, i.e., the functions satisfy $\sum_{j=1}^{d-1}l_{j}(x)=1$ and $\sum_{j=1}^{d-1}l_{j}(x)\cdot\pi_{w^{\perp}}(b_{1}-b_{j})=x$ for every $x\in M$. For any $1\geq\alpha\geq 0$, observe that $l_{1}$ takes values in the interval $\alpha,1]$ on the set $(1-\alpha)M$. Hence we get
@@
$\geq(1-\alpha)^{d-2}\max_{x\in M}e^{-dL\|\alpha x\|},$
where in the last inequality, we use $dL$-log-Lipschitzness of $\mu^{\prime}$ to see that for any $s\in M$ we have $\frac{\mu^{\prime}((1-\alpha)s)}{\mu^{\prime}(s)}\leq\max_{x\in M}e^{-dL\|\alpha x\|}$.
@@
where the last inequality comes from $d\geq 3$ and $L\geq 1$. Thus (22) holds as desired.
5.7 Proof of Lemma 5.20
To show Lemma 5.20 on the width of the facet, we need the following upper bound about the mass around zero for a random variable whose density is formed by a log-Lipschitz function multiplied by a convex function. Its function is to deal with the volume term that we receive from the Jacobian in Lemma 2.13.
Lemma 5.25.